Shalimar Foods sells Vegetarian Corn Samosa plates for INR 30 each, and serves an average of 625

customers per week. During a recent promotion on the eve of India’s Independence Day

, Shalimarcut its price to INR 25 each and observed an increase in sales to 750 plates per week.a.

 

Calculate Shalimar’s price elasticity of Demand

 b.

 

Shalimar is considering permanently lowering its price to INR 28 to increase revenue. Explainc.

 

How many plates should Shalimar expect to sell at the new price?d.

 

Does the move make sense in the light of Shalimar’s intent to increase revenue?


1
Expert's answer
2021-11-16T10:40:36-0500

Given:

Old price=$30

Old quantity=625

New price=$25

New quantity=750


a. The price elasticity of Demand:

= % change in quantity / % change in price

=(new quantity−old quantity/old quantity]*100

/(new price−old price/old price*100)

"=\frac{ [\frac{(750 − 625) }{ 625}]\times 100}{[\frac{(25 − 30) }{ 30}]\times100 }\\

= \frac{20\% }{ − 16.67\%} \\

= − 1.2"

Hence, The price elasticity of demand is -1.2.

b. As the elasticity is more than 1, the demand is elastic. One percent decrease or increase in price would lead to a decrease or increase in the quantity being demanded by more than one percent. This would lead to an increase or decrease in revenue.

Hence, increasing the price to Rs 28 from Rs 25 would be an incorrect decision.

c. New quantity:

Elasticity= % change in quantity "\div"  % change in price

"−1.2= \frac{[\frac{(Q − 625) }{ 625}]\times100}{[\frac{(25 − 30) }{ 30}]\times100} \\


−20.004\times 625=100Q−62500\\

=500"


Hence, the new quantity would be 500.

d. The decision of increasing price from Rs 25 to Rs 28 would be considered to be unwise as revenue would tend to decrease from its present level.